OpenStudy (anonymous):

let g(x)=lim h->0 (f(x+h)-f(x))/h for f(x)=5/(x+5)^1/2 Find g(4)

OpenStudy (unklerhaukus):

\[g(x) = f'(x) = -5/2(x+5)^{3/2} \] \[g(4)= f'(4) = -5/2(4+5^{3/2})=~-166.25\]

OpenStudy (unklerhaukus):

i mean ~ -0.0376